# Gaussian Process-Based Prediction and Control of Hammerstein-Wiener Systems

### v1

@@ IV Data-Driven Predictive Control with Implicit Gaussian Process Model
where $c_{p}:=M\left(\sqrt{\mu(p)\sigma_{\text{p}}(\eta^{k})}+\left\lVert m_{\text{p}}(\eta^{k})\right\rVert_{2}\right)$, $\sigma_{\text{p}}(\eta^{k})$ is the largest eigenvalue of $\hat{k}_{\text{p}}(\eta^{k})$ and $\mu(\cdot)$ is the quantile function of the $\chi^{2}$-distribution with $n_{y}L^{\prime}$ degrees of freedom.
Proof:
We have
$\mathrm{Pr}\left(\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})-m_{\text{p}}(\eta^{k})\right\rVert^{2}_{\hat{k}_{\text{p}}^{-1}(\eta^{k})}\leq\mu(p)\right)=p,$
since $\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})-m_{\text{p}}(\eta^{k})\right\rVert^{2}_{\hat{k}_{\text{p}}^{-1}(\eta^{k})}$ follows a $\chi^{2}$-distribution with $n_{y}L^{\prime}$ degrees of freedom. According to Assumptions 1.3) and 2, we have
$\begin{split}&\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})-m_{\text{p}}(\eta^{k})\right\rVert_{\hat{k}_{\text{p}}^{-1}(\eta^{k})}\\ \geq&\ \tfrac{1}{\sqrt{\sigma_{\text{p}}(\eta^{k})}}\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})-m_{\text{p}}(\eta^{k})\right\rVert_{2}\\ \geq&\ \tfrac{1}{\sqrt{\sigma_{\text{p}}(\eta^{k})}}\left(\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})\right\rVert_{2}-\left\lVert m_{\text{p}}(\eta^{k})\right\rVert_{2}\right)\\ \geq&\ \tfrac{1}{M\sqrt{\sigma_{\text{p}}(\eta^{k})}}\left\lVert\mathbf{y}^{k}_{0}-\mathbf{y}^{k}\right\rVert_{2}^{2}-\tfrac{1}{\sqrt{\sigma_{\text{p}}(\eta^{k})}}\left\lVert m_{\text{p}}(\eta^{k})\right\rVert_{2}.\end{split}$
This leads to $\mathrm{Pr}\left(\left\lVert\mathbf{y}^{k}_{0}-\mathbf{y}^{k}\right\rVert_{2}\leq c_{p}\right)\geq p$. If $\left\lVert\mathbf{y}^{k}_{0}-\mathbf{y}^{k}\right\rVert_{2}\leq c_{p}$, according to [33, Theorem 2.3], $H\mathbf{y}^{k}_{0}\leq q$ is guaranteed by
$h_{i}^{\top}\mathbf{y}^{k}\leq q_{i}-\tau_{V}(h_{i}),\quad i=1,\dots,n_{c},$ (36)
where $H=:\mathrm{col}\left(h_{1}^{\top},\dots,h_{n_{c}}^{\top}\right)$, $q=:\mathrm{col}\left(q_{1},\dots,q_{n_{c}}\right)$, and $\tau_{V}(h_{i})=c_{p}\sqrt{h_{i}^{\top}h_{i}}$ denotes the support function of set $V:=\left\{\mathbf{e}\,|\left\lVert\mathbf{e}\right\rVert_{2}\leq c_{p}\right\}$. Aggregating (36) for all $i$ directly leads to the lemma. ∎
From the distribution (34), the expected control cost can be calculated as
$J_{\text{ctr}}=\left\lVert\mathbf{u}^{k}\right\rVert_{R}^{2}+\left\lVert\Phi(\mathbf{y}^{k})+m_{\text{p}}(\eta^{k})-\Phi(\mathbf{r}^{k})\right\rVert_{Q}^{2}+\mathrm{tr}\left(Q\hat{k}_{\text{p}}(\eta^{k})\right).$
Instead of providing an output prediction for a particular input, the control problem optimizes $\mathbf{u}^{k}$ and $\mathbf{y}^{k}$ at the same time by penalizing a combination of the nominal tracking error $\left\lVert\Phi(\mathbf{y}^{k})+m_{\text{p}}(\eta^{k})-\Phi(\mathbf{r}^{k})\right\rVert_{Q}^{2}$ and the prediction error $\mathrm{tr}\left(Q\hat{k}_{\text{p}}(\eta^{k})\right)$. This avoids applying Algorithm 1 directly in the control problem, i.e., solving (24), (25), or (26) as an inner problem. Since the output nonlinearity $\Phi(\cdot)$ is not known exactly, $\Phi(\mathbf{y}^{k})$ and $\Phi(\mathbf{r}^{k})$ are replaced by their posterior means $\mathbf{m}_{y\text{p}}(\mathbf{y}^{k})$ and $\mathbf{m}_{y\text{p}}(\mathbf{r}^{k})$ with certainty equivalence.
@@
$\underset{\mathbf{u}^{k},\mathbf{y}^{k}}{\text{min}}&\left\lVert\mathbf{u}^{k}\right\rVert_{R}^{2}+\left\lVert\mathbf{m}_{y\text{p}}(\mathbf{y}^{k})+m_{\text{p}}(\eta^{k})-\mathbf{m}_{y\text{p}}(\mathbf{r}^{k})\right\rVert_{Q}^{2}+\mathrm{tr}\left(Q\hat{k}_{\text{p}}(\eta^{k})\right)\\ \text{s.t. }&\ \mathbf{u}^{k}\in\mathcal{U}_{k},\ H\mathbf{y}^{k}\leq q-c_{p}\sqrt{\mathrm{diag}\left(HH^{\top}\right)}.\end{aligned}$
Then, the first entry in the optimal input sequence is applied to the system and the noisy output $y_{k}$ is measured.
V Numerical Examples
@@
Consider Hammerstein-Wiener systems with input and output nonlinearities $\psi(u)=u+\sin(u)$, $\phi^{-1}(\bar{y})=\bar{y}+\sin(\bar{y})$ and random single-input, single-output second-order systems for the linear part $G(q)$, generated by the Matlab function drss and normalized to have an $\mathcal{H}_{2}$-norm of 10. The output nonlinearity satisfies Assumption 2 with $M=2$. The following parameters are used: $L_{0}=2$, $N=100$, $\sigma=0.01$. Unit Gaussian inputs are used for both training and testing. The hyper-hyperparameters $\zeta=[\lambda\ \alpha]^{\top}$ are selected by cross-validation with grid $\lambda\in\{0.25,0.5,1,2,4\}$, $\alpha\in\{0.5,0.6,0.7,0.8,0.9\}$. The squared exponential kernel is used for $k_{u}(\cdot,\cdot)$ and $k_{y}(\cdot,\cdot)$, i.e., $k_{u}(u^{\prime},u^{\prime\prime};l_{u})=\exp\left(-\frac{(u^{\prime}-u^{\prime\prime})^{2}}{2l_{u}^{2}}\right)$ with hyperparameter $l_{u}$, and similarly for $k_{y}(y^{\prime},y^{\prime\prime};l_{y})$. Note that the scaling factor is omitted as discussed in Remark 4. The prior mean functions are selected as identity, i.e., $m_{u}(u)=u$, $m_{y}(y)=y$. The minimum mean-squared error objective (25) is used to obtain explicit predictions in Algorithm 1 and solved using the Matlab function particleswarm with a particle swarm solver. The other optimization problems (29) and (37) are solved using the Matlab function fmincon with the sequential quadratic programming algorithm.
The proposed implicit Gaussian process model is benchmarked against a black-box Gaussian process model of
$y_{\text{f},l}\sim\mathcal{N}\left(m_{\text{bb},l}(\mathrm{col}\left(\mathbf{u},\mathbf{y}_{\text{p}}\right)),\sigma_{\text{bb},l}^{2}(\mathrm{col}\left(\mathbf{u},\mathbf{y}_{\text{p}}\right))\right),$
where $l=1,\dots,L^{\prime}$ and $m_{\text{bb},l}(\cdot)$, $\sigma_{\text{bb},l}(\cdot)$ are the posterior output prediction and standard deviation of the $l$-step-ahead predictor, respectively. The model is trained using default settings of the Matlab function fitrgp. We also compare a linear predictor based on Willems’ fundamental lemma, neglecting the nonlinearities: $\mathbf{y}_{\text{f}}=H_{y\text{f}}\,\mathrm{col}\left(H_{u},H_{y\text{p}}\right)^{\dagger}\mathrm{col}\left(\mathbf{u},\mathbf{y}_{\text{p}}\right)$, where $H_{y}=:\mathrm{col}\left(H_{y\text{p}},H_{y\text{f}}\right)$, $H_{y\text{p}}\in\mathbb{R}^{n_{y}L_{0}\times M}$, and $H_{y\text{f}}\in\mathbb{R}^{n_{y}L^{\prime}\times M}$.
@@
For demonstrative purposes, we first consider the case where $L^{\prime}=1$. Figure 2 demonstrates the prediction accuracy of the predictors on a test data set. The recovered input and output nonlinear functions are plotted in Figure 3 by applying (32) to a linear grid of query points between $-\pi$ and $\pi$. It can be seen from Figure 2 that Algorithm 1 obtains very accurate predictions and performs much better than the black-box Gaussian process model and the linear predictor. The black-box predictions are similar to the linear ones, indicating that the black-box Gaussian process model does not capture the nonlinearities well. As can be seen from Figure 3, the shapes of the nonlinearities are captured correctly. There is a discrepancy of $\psi(u)$ in terms of the scaling, which is expected since only the normalized nonlinearities are recovered.
@@
Figure 2: Comparison of data-driven prediction with different predictors.
@@
Figure 3: Estimation of input and output nonlinear functions $\psi(u)$ and $\phi(y)$.
@@
Then, multi-step-ahead predictors with $L^{\prime}=4$ are considered by conducting 50 Monte Carlo simulations. At each time $k$, $l$-step-ahead prediction accuracy is evaluated quantitatively on a test data set by comparing the $l$-step-ahead prediction at time $k$ with the true output at time $(k+l)$, $l=1,2,3,4$. In addition to the predictors presented before, we also consider the implicit Gaussian process model with hyperparameters estimated without the hyperprior, i.e., $S_{\gamma}^{-1}=\mathbf{0}$ in (29). The median root-mean-squared prediction errors are shown in Table I. Similar to one-step-ahead prediction, our proposed algorithm achieves the smallest prediction errors across the horizon, with an overall error reduction of 40% and 38% compared to the black-box Gaussian process model and the linear predictor, respectively. On the other hand, a significant prediction error is observed for Algorithm 1 without using the hyperprior in hyperparameter estimation. This indicates that the proposed method would fail without the hyperprior and validates the effectiveness of the proposed JMAP-ML scheme for hyperparameter estimation.
@@
TABLE I: Median root-mean-squared prediction errors of different predictors with $L^{\prime}=4$
$l$ 1 2 3 4
Algorithm 1 3.532 3.742 4.171 3.983
Algorithm 1 w/o hyperprior 17.074 17.101 17.281 18.067
Black-box model 6.062 6.439 6.528 6.708
Linear predictor 5.965 6.317 6.395 6.398
Finally, the data-driven predictive control algorithm (37) is applied where we consider the following second-order system for $G(q)$:
$G(q)=\frac{10q}{q^{2}+0.24q+0.36}$
with prediction horizon $L^{\prime}=4$. We have also implemented receding horizon control for the black-box and the linear predictors. For the black-box Gaussian process model, the following optimization problem is solved at time $k$:
$\underset{\mathbf{u}^{k},\mathbf{y}^{k}}{\text{min}}$ $\ \left\lVert\mathbf{u}^{k}\right\rVert_{R}^{2}+\left\lVert\mathbf{y}^{k}-\mathbf{r}^{k}\right\rVert_{Q}^{2}+\mathrm{diag}\left(Q\right)^{\top}\boldsymbol{\sigma}^{2}_{\text{bb}}(\xi^{k})$
s.t. $\ \mathbf{y}^{k}=\mathbf{m}_{\text{bb}}(\xi^{k}),\ \mathbf{u}^{k}\in\mathcal{U}_{k},$
$\ H\mathbf{y}^{k}\leq q-\bar{\sigma}_{\text{bb}}(\xi^{k})\sqrt{\mu(p)\mathrm{diag}\left(HH^{\top}\right)},$
where $\xi^{k}=\mathrm{col}\left(u_{k-L_{0}},\dots,u_{k-1},\mathbf{u}^{k},y_{k-L_{0}},\dots,y_{k-1}\right)$ and $\bar{\sigma}_{\text{bb}}(\xi^{k})$ denotes the largest element of $\boldsymbol{\sigma}_{\text{bb}}(\xi^{k})$. The derivation is similar to Section IV. For the linear predictor, we apply subspace predictive control [28] with certainty equivalence.

### v2

@@ 4 Data-Driven Predictive Control with Implicit Gaussian Process Model
where $c_{p}:=M\left(\sqrt{\mu_{\chi^{2}}(p)\sigma_{\mathrm{p}}(\eta^{k})}+\left\lVert m_{\mathrm{pm}}(\eta^{k})\right\rVert_{2}\right)$, $\sigma_{\mathrm{p}}(\eta^{k})$ is the largest eigenvalue of $\hat{k}_{\mathrm{pm}}(\eta^{k})$ and $\mu_{\chi^{2}}(\cdot)$ is the quantile function of the $\chi^{2}$-distribution with $n_{y}L^{\prime}$ degrees of freedom.
{pf}
From (53), we have
$\mathrm{Pr}\left(\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})-m_{\mathrm{pm}}(\eta^{k})\right\rVert^{2}_{\hat{k}_{\mathrm{pm}}^{-1}(\eta^{k})}\leq\mu_{\chi^{2}}(p)\right)=p,$
since $\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})-m_{\mathrm{pm}}(\eta^{k})\right\rVert^{2}_{\hat{k}_{\mathrm{pm}}^{-1}(\eta^{k})}$ follows a $\chi^{2}$-distribution with $n_{y}L^{\prime}$ degrees of freedom. Using Assumption 2 and the reverse triangle inequality, we have
$\begin{split}&\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})-m_{\mathrm{pm}}(\eta^{k})\right\rVert_{\hat{k}_{\mathrm{pm}}^{-1}(\eta^{k})}\\ \geq&\ \tfrac{1}{\sqrt{\sigma_{\mathrm{p}}(\eta^{k})}}\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})-m_{\mathrm{pm}}(\eta^{k})\right\rVert_{2}\\ \geq&\ \tfrac{1}{\sqrt{\sigma_{\mathrm{p}}(\eta^{k})}}\left(\left\lVert\Phi(\mathbf{y}^{k}_{0})-\Phi(\mathbf{y}^{k})\right\rVert_{2}-\left\lVert m_{\mathrm{pm}}(\eta^{k})\right\rVert_{2}\right)\\ \geq&\ \tfrac{1}{M\sqrt{\sigma_{\mathrm{p}}(\eta^{k})}}\left\lVert\mathbf{y}^{k}_{0}-\mathbf{y}^{k}\right\rVert_{2}-\tfrac{1}{\sqrt{\sigma_{\mathrm{p}}(\eta^{k})}}\left\lVert m_{\mathrm{pm}}(\eta^{k})\right\rVert_{2}.\end{split}$
This leads to $\mathrm{Pr}\left(\left\lVert\mathbf{y}^{k}_{0}-\mathbf{y}^{k}\right\rVert_{2}\leq c_{p}\right)\geq p$. If $\left\lVert\mathbf{y}^{k}_{0}-\mathbf{y}^{k}\right\rVert_{2}\leq c_{p}$, according to [21, Theorem 2.3], $H\mathbf{y}^{k}_{0}\leq q$ is guaranteed by
$h_{i}^{\top}\mathbf{y}^{k}\leq q_{i}-\rho_{V}(h_{i}),\quad i=1,\dots,n_{c},$ (55)
where $H=:\mathrm{col}\left(h_{1}^{\top},\dots,h_{n_{c}}^{\top}\right)$, $q=:\mathrm{col}\left(q_{1},\dots,q_{n_{c}}\right)$, and $\rho_{V}(h_{i})=c_{p}\sqrt{h_{i}^{\top}h_{i}}$ denotes the support function of the set $V:=\left\{\mathbf{e}\,|\left\lVert\mathbf{e}\right\rVert_{2}\leq c_{p}\right\}$. Aggregating (55) for all $i$ directly leads to the lemma.∎
@@
$\underset{\mathbf{u}^{k},\mathbf{y}^{k}}{\mathrm{min}}$ $\left\lVert\mathbf{u}^{k}\right\rVert_{R}^{2}+\left\lVert\mathbf{m}_{y\mathrm{p}}(\mathbf{y}^{k})+m_{\mathrm{pm}}(\eta^{k})-\mathbf{m}_{y\mathrm{p}}(\mathbf{r}^{k})\right\rVert_{Q}^{2}$ (56)
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\ \ +\mathrm{tr}\left(Q\hat{k}_{\mathrm{pm}}(\eta^{k})\right)$
$\mathrm{s.t.}$ $\ \mathbf{u}^{k}\in\mathcal{U}_{k},\ H\mathbf{y}^{k}\leq q-c_{p}\sqrt{\mathrm{diag}\left(HH^{\top}\right)},$
where $\mathbf{m}_{y\mathrm{p}}(\cdot)$ is the estimated output nonlinearity given by (49), $m_{\mathrm{pm}}(\cdot)$ is the posterior mean of the implicit function given by (33), and $\hat{k}_{\mathrm{pm}}(\cdot)$ is the modified posterior covariance of the implicit function given by (34) and (39). Then, the first entry in the optimal input sequence is applied to the system, and the noisy output $y_{k}$ is measured.
As discussed above, there is no explicit predictor in (56). Instead, the implicit predictor is embedded in the expected cost function and the constraint tightening term, penalizing both the bias ($m_{\mathrm{pm}}(\eta^{k})$) and the variance ($\hat{k}_{\mathrm{pm}}(\eta^{k})$) errors. This implicitly prohibits unreliable choices of $\eta^{k}$, which are unlikely to be generated by the system.
5 Numerical Examples
@@
The following parameters are used throughout this section: $L_{0}=n_{x}$, $N=100$, $\sigma=0.01$. The hyper-hyperparameters $\zeta=[\lambda\ \alpha]^{\top}$ are selected by cross-validation with grid $\lambda\in\{0.25,0.5,1,2,4\}$, $\alpha\in\{0.5,0.6,0.7,0.8,0.9\}$. The squared exponential kernel is used for $k_{u}(\cdot,\cdot)$ and $k_{y}(\cdot,\cdot)$, i.e., $k_{u}(u^{\prime},u^{\prime\prime};l_{u})=\exp\left(-\frac{(u^{\prime}-u^{\prime\prime})^{2}}{2l_{u}^{2}}\right)$ with hyperparameter $l_{u}$, and similarly for $k_{y}(y^{\prime},y^{\prime\prime};l_{y})$. Note that the scaling factor is omitted as discussed in Remark 3. The prior mean functions are selected as identity, i.e., $m_{u}(u)=u$, $m_{y}(y)=y$. The minimum mean-squared error objective (42) is used to obtain explicit predictions in Algorithm 2 and solved using the Matlab function particleswarm with a particle swarm solver. The other optimization problems (46) and (56) are solved using the Matlab function fmincon with the interior point algorithm.
The proposed implicit GP model is benchmarked against a black-box GP model without incorporating the H-W model structure. To obtain multi-step-ahead predictions, denoted by
$y_{\mathrm{f},l}\sim\mathcal{N}\left(m_{\mathrm{bb},l}(\mathrm{col}\left(\mathbf{u},\mathbf{y}_{\mathrm{p}}\right)),\sigma_{\mathrm{bb},l}^{2}(\mathrm{col}\left(\mathbf{u},\mathbf{y}_{\mathrm{p}}\right))\right),$
where $m_{\mathrm{bb},l}(\cdot)$, $\sigma_{\mathrm{bb},l}(\cdot)$ are the posterior nominal prediction and standard deviation of the $l$-step-ahead predictor, respectively, two strategies are considered. 1) Black-box posterior predictors for all prediction horizons $l=1,\dots,L^{\prime}$ are learned separately. 2) The one-step-ahead black-box predictor is applied recursively with nominal predictions propagated with certainty equivalence. The models are trained using the Matlab function fitrgp with hyperparameters estimated by bayesopt using 100 evaluations.
We also compare a linear predictor based on WFL, neglecting the nonlinearities:
$\mathbf{y}_{\mathrm{f}}=H_{y\mathrm{f}}\,\mathrm{col}\left(H_{u},H_{y\mathrm{p}}\right)^{\dagger}\mathrm{col}\left(\mathbf{u},\mathbf{y}_{\mathrm{p}}\right),$
where $H_{y}=:\mathrm{col}\left(H_{y\mathrm{p}},H_{y\mathrm{f}}\right)$, $H_{y\mathrm{p}}\in\mathbb{R}^{n_{y}L_{0}\times M}$, and $H_{y\mathrm{f}}\in\mathbb{R}^{n_{y}L^{\prime}\times M}$.
5.1 Data-Driven Prediction
We first benchmark the performance of Algorithm 2. In this subsection, the linear part of the model $G(q)$ is selected as a random single-input, single-output second-order system, generated by the Matlab function drss. The systems are normalized to have an $\mathcal{H}_{2}$-norm of 10. Gaussian inputs with a standard deviation of 2 are used for both training and testing.
For demonstrative purposes, we first investigate one-step-ahead prediction with $L^{\prime}=1$ and fixed input and output nonlinearities $\psi(u)=u+\sin(u)$, $\phi^{-1}(\bar{y})=\bar{y}+\sin(\bar{y})$. We consider two cases: 1) with 20 virtual derivative points in a linear grid between $\pi$ and $\pi$ and 2) without virtual derivative points $n_{m}=0$, indicated by “w/o mono.” in the figures.
@@ 5.1 Data-Driven Prediction
Figure 2 demonstrates the prediction accuracy of the predictors on a 50-point test dataset. The recovered input and output nonlinear functions are plotted in Figure 3 by applying (48)-(51) to a linear grid of query points between $\pi$ and $\pi$. It can be seen from Figure 2 that Algorithm 2 obtains very accurate predictions and performs much better than the black-box GP model and the linear predictor. The addition of virtual derivative points does not have a large effect on the predictions. The black-box predictions are similar to the linear ones, indicating that the black-box GP model does not capture the nonlinearities well. As can be seen from Figure 3, the shapes of the nonlinearities are captured accurately by Algorithm 2 with virtual derivative points, whereas the estimated output nonlinearity is not monotonic when virtual derivative points are absent. This proves the effectiveness of promoting monotonicity by introducing virtual derivative points. There is a discrepancy of $\psi(u)$ in terms of the scaling, which is expected since only the normalized nonlinearities are recovered, cf. Remark 3.
@@
Figure 2: Comparison of data-driven prediction with different predictors. W/o mono.: no virtual derivative points.
@@
Figure 3: Estimation of input and output nonlinear functions $\psi(u)$ and $\phi(y)$. W/o mono.: no virtual derivative points.
Then, multi-step-ahead predictors with $L^{\prime}=4$ are considered by conducting 100 Monte Carlo simulations. The configurations are the same as the one-step-ahead case, except that random input and output nonlinearities are considered by generating samples from the following Gaussian process: $f(\cdot)\sim\mathcal{GP}\left(m(\cdot),k(\cdot,\cdot)\right)$, $m(\eta)=\eta$, $k(\eta^{\prime},\eta^{\prime\prime})=\exp\left(-(\eta^{\prime}-\eta^{\prime\prime})^{2}/8\right)$, and the virtual derivative point are selected between -5 and 5. The output nonlinearity is verified to be monotonically increasing.
@@
At each time $k$, $l$-step-ahead prediction accuracy is evaluated quantitatively on a 50-point test data set by comparing the $l$-step-ahead prediction at time $k$ with the true output at time $(k+l)$, $l=1,2,3,4$. The boxplot of root-mean-squared prediction errors is shown in Figure 4. In addition to the predictors presented before, we also consider the implicit GP model with hyperparameters estimated without the hyperprior, i.e., $S_{\gamma}^{-1}=\mathbf{0}$ in (46). This is indicated by “w/o hyperprior” in Figure 4.
Similar to one-step-ahead prediction, our proposed algorithm achieves the smallest prediction errors over the whole horizon, with overall median error reductions of 59.8%, 58.8%, and 70.5% compared to the multi-step black-box GP model, the recursive one-step GP model, and the linear predictor, respectively. On the other hand, a significant prediction error is observed when the hyperprior is not used in hyperparameter estimation. This indicates that the proposed method would fail without the hyperprior, validating the effectiveness of the proposed JMAP-ML scheme for hyperparameter estimation.
Admittedly, our proposed approach suffers from high computational complexity, with a median training time of 69 s for each hyper-hyperparameter choice and a median prediction time of 56 s for 50 predictions on an AMD Ryzen 7 8845H laptop. In comparison, the black-box GP model requires a median training time of 46 s and a median prediction time of 0.02 s.
@@
Figure 4: Boxplot of root-mean-squared prediction errors for different predictors with $L^{\prime}=4$. $l$: step of forward prediction.
5.2 Data-Driven Predictive Control
Finally, the DDPC algorithm (56) is applied with prediction horizon $L^{\prime}=4$. In this subsection, we consider the following pH process model obtained in [1]:
$G(q)$ $=\frac{-1.728q^{3}+q^{2}+0.5056q}{q^{3}-0.797q^{2}-0.1609q+0.06203},$
$\psi(u)$ $=\frac{1-e^{-u}}{1+e^{-u}},\quad\phi^{-1}(\bar{y})=\frac{1-e^{-\bar{y}}}{1+e^{-\bar{y}}}.$